OpenStudy (anonymous):

A small applience manufacturer finds that it costs $9000 to make 1000 toasters per week and $12,000 to make 1500 per week Express the cost as a function of the number of toasters produced what is the slope of the graph and what does it represent whats the y intercept

OpenStudy (anonymous):

what is the line through the points (1000,9000) and (1500,12000) The slope represents how much it costs to make one appliance the y-intercept represents how much it costs even though the appliance manufacturer doesn't make anything

OpenStudy (anonymous):

y=6x+3000

OpenStudy (anonymous):

do you know what the expressed function is ?

OpenStudy (anonymous):

ah how do you find that?

OpenStudy (anonymous):

well, what is the line passing through (1000,9000) and (1500,12000) So, the slope of the line is m=(y1-y2)/(x1-x2) m=6 Then choose one point and plug into point-slope form y-9000=6(x-1000) y-9000=6x-6000 y=6x+3000

OpenStudy (anonymous):

that makes so much more sense thank you for your time